# Residual-based Chebyshev filtered subspace iteration for sparse Hermitian eigenvalue problems tolerant to inexact matrix-vector products

### v2

@@ 4.1.2 Hermitian Eigenvalue Problems
We now consider the finite-element discretization of the Kohn-Sham DFT equations at a non-zero $k$-point  [38] in the Brillouin zone for the systems described in Table 1. The resulting discretized equation is now a complex Hermitian eigenvalue problem. The computation of the matrix multi-vector products with FP64 arithmetic is done according to the methodology prescribed by [38, 55]. When using FP32 arithmetic to evaluate $\textbf{A}\textbf{R}_{\textbf{Y}}$ in Algorithm 2 of Algorithm 2, the data structures storing the matrix A and multi-vectors $\textbf{R}_{\textbf{X}}$ and $\textbf{R}_{\textbf{Y}}$ are in FP32 format and the BLAS calls used to perform $\textbf{A}\textbf{R}_{\textbf{Y}}$ are replaced with corresponding FP32 variants while the addition and scaling of multi-vectors is done using FP64 arithmetic. For TF32 arithmetic on NVIDIA GPUs we replace the cublasCgemm and its strided/batched variants with cublasGemmEx and its strided/batched variants with computeType set to CUBLAS_COMPUTE_32F_FAST_TF32.
@@ 4.2 Generalized Eigenvalue Problems
Input:
Chebyshev polynomial order $p$, estimates of the bounds of the eigenspectrum $\lambda_{max},\lambda_{min}$, estimate of the upper bound of the wanted spectrum $\lambda_{T}$ and the initial guess of eigenvectors $\textbf{X}^{(i)}$ and eigenvalues $\bm{\Lambda}^{(i)}$
Result: The filtered subspace $\textbf{Y}_{p}^{(i)}$
Temporary Variable: X, Y, $\textbf{R}_{\textbf{X}}$, $\textbf{R}_{\textbf{Y}}$,$\bm{\Lambda}_{\textbf{X}}$ and $\bm{\Lambda}_{\textbf{Y}}$
1 $e\leftarrow\frac{\lambda_{max}-\lambda_{T}}{2}$; $c\leftarrow\frac{\lambda_{max}+\lambda_{T}}{2}$;
2 $\sigma\leftarrow\frac{e}{\lambda_{min}-c}$;
3 $\sigma_{1}\leftarrow\sigma$;
4 $\gamma\leftarrow\frac{2}{\sigma_{1}}$;
5 $\textbf{X}\leftarrow\textbf{X}^{(i)}$;
6 $\textbf{Y}\leftarrow\textbf{A}\textbf{X}^{(i)}-\textbf{B}\textbf{X}^{(i)}\bm{\Lambda}^{(i)}$;
7 $\textbf{R}_{\textbf{X}}\leftarrow 0$;
8 $\textbf{R}_{\textbf{Y}}\leftarrow\frac{2\sigma_{1}}{e}\textbf{Y}$;
9 $\bm{\Lambda}_{\textbf{X}}\leftarrow\textbf{I}$;
10 $\bm{\Lambda}_{\textbf{Y}}\leftarrow\frac{2\sigma_{1}}{e}\left(\bm{\Lambda}^{(i)}-c\textbf{I}\right)$;
11 for $k\leftarrow 2$ to $p$ do
12       $\sigma_{2}\leftarrow\frac{1}{\gamma-\sigma}$;
13       $\textbf{R}_{\textbf{X}}\leftarrow\frac{2\sigma_{2}}{e}\textbf{A}\textbf{D}^{-1}\textbf{R}_{\textbf{Y}}-\frac{2\sigma_{2}}{e}c\textbf{R}_{\textbf{Y}}-\sigma\sigma_{2}\textbf{R}_{\textbf{X}}+\frac{2\sigma_{2}}{e}\textbf{Y}\bm{\Lambda}_{\textbf{Y}}$;
14       $\bm{\Lambda}_{\textbf{X}}\leftarrow\frac{2\sigma_{2}}{e}\bm{\Lambda}_{\textbf{Y}}\bm{\Lambda}^{(i)}-\frac{2\sigma_{2}}{e}c\bm{\Lambda}_{\textbf{Y}}-\sigma\sigma_{2}\bm{\Lambda}_{\textbf{X}}$;
@@ 12       $\sigma_{2}\leftarrow\frac{1}{\gamma-\sigma}$;
16       swap$\left(\bm{\Lambda}_{\textbf{X}},\bm{\Lambda}_{\textbf{Y}}\right)$;
17       $\sigma=\sigma_{2}$;
18
[…]
@@ 17       $\sigma=\sigma_{2}$;
Algorithm 3 Reformulated ChFSI (R-ChFSI) Procedure for generalized eigenvalue problems
@@ 4.2.1 Symmetric Eigenvalue Problems
From Fig. 7, we observe that for all the benchmark systems summarized in Table 1, the residual tolerance than can be achieved by the R-ChFSI method is nearly 10 orders of magnitude lower than what can be achieved using the ChFSI method when employing the diagonal approximation for the inverse overlap matrix. We also note that our proposed R-ChFSI algorithm allows us to construct the filtered subspace using lower precision and we report the residuals achieved by the FP32 and TF32 variants of R-ChFSI in Fig. 8. We also report the speedups achieved by the FP32 and TF32 variants of R-ChFSI over the FP64 variant of R-ChFSI in Fig. 9 on NVIDIA H100 GPU accelerators. We achieve speedups of up to 2.3x for the TF32 variant of the R-ChFSI method over the baseline implementation of the ChFSI method with FP64 arithmetic.
@@ 4.2.2 Hermitian Eigenvalue Problems
We now consider the finite-element discretization of the Kohn-Sham DFT equations at a non-zero $k$-point  [38] in the Brillouin zone for the systems described in Table 1. The resulting discretized equation is now a complex Hermitian eigenvalue problem. All the computations are done with FP64 according to the methodology prescribed by [38, 55]. From Fig. 10, we observe that for all the benchmark systems summarized in Table 1, even for Hermitian eigenvalue problems the residual tolerance than can be achieved by the R-ChFSI method is nearly 10 orders of magnitude lower than what can be achieved using the ChFSI method when employing the diagonal approximation for the inverse overlap matrix. We also note that our proposed R-ChFSI algorithm allows for us to construct the filtered subspace using lower precision and we report the residuals achieved by the FP32 and TF32 variants of R-ChFSI in Fig. 11. We also report the speedups achieved by the FP32 and TF32 variants of R-ChFSI over the FP64 variant of R-ChFSI in Fig. 12 on NVIDIA H100 GPU accelerators. We achieve speedups of up to 2.7x for the TF32 variant of the R-ChFSI method over the baseline implementation of the R-ChFSI method with FP64 arithmetic.

### v3

@@ 4.1.2 Hermitian Eigenvalue Problems
We now consider the finite-element discretization of the Kohn-Sham DFT equations at a non-zero $k$-point  [7] in the Brillouin zone for the systems described in Table 1. The resulting discretized equation is now a complex Hermitian eigenvalue problem. The computation of the matrix multi-vector products with FP64 arithmetic is done according to the methodology prescribed by [7, 28]. When using FP32 arithmetic to evaluate $\textbf{A}\textbf{R}_{\textbf{Y}}$ in Algorithm 2, the data structures storing the matrix A and multi-vectors $\textbf{R}_{\textbf{X}}$ and $\textbf{R}_{\textbf{Y}}$ are in FP32 format and the BLAS calls used to perform $\textbf{A}\textbf{R}_{\textbf{Y}}$ are replaced with corresponding FP32 variants while the addition and scaling of multi-vectors is done using FP64 arithmetic. For TF32 arithmetic on NVIDIA GPUs we replace the cublasCgemm and its strided/batched variants with cublasGemmEx and its strided/batched variants with computeType set to CUBLAS_COMPUTE_32F_FAST_TF32.
@@ 4.2 Generalized Eigenvalue Problems
Algorithm 3 R-ChFSI procedure for generalized Hermitian eigenvalue problems
INPUTS: Chebyshev polynomial order $p$, estimates of the bounds of the eigenspectrum $\lambda_{max},\lambda_{min}$, estimate of the upper bound of the wanted spectrum $\lambda_{T}$ and the initial guess of eigenvectors $\textbf{X}^{(i)}$ and eigenvalues $\boldsymbol{\Lambda}^{(i)}$
OUTPUT: The filtered subspace $\textbf{Y}_{p}^{(i)}$
TEMPORARY VARIABLES:X, Y, $\textbf{R}_{\textbf{X}}$, $\textbf{R}_{\textbf{Y}}$,$\boldsymbol{\Lambda}_{\textbf{X}}$ and $\boldsymbol{\Lambda}_{\textbf{Y}}$
  $e\leftarrow\frac{\lambda_{max}-\lambda_{T}}{2}$; $c\leftarrow\frac{\lambda_{max}+\lambda_{T}}{2}$; $\sigma\leftarrow\frac{e}{\lambda_{min}-c}$; $\sigma_{1}\leftarrow\sigma$; $\gamma\leftarrow\frac{2}{\sigma_{1}}$
  $\textbf{X}\leftarrow\textbf{X}^{(i)}$; $\textbf{Y}\leftarrow\textbf{A}\textbf{X}^{(i)}-\textbf{B}\textbf{X}^{(i)}\boldsymbol{\Lambda}^{(i)}$
  $\textbf{R}_{\textbf{X}}\leftarrow 0$; $\textbf{R}_{\textbf{Y}}\leftarrow\frac{2\sigma_{1}}{e}\textbf{Y}$
  $\boldsymbol{\Lambda}_{\textbf{X}}\leftarrow\textbf{I}$; $\boldsymbol{\Lambda}_{\textbf{Y}}\leftarrow\frac{2\sigma_{1}}{e}\left(\boldsymbol{\Lambda}^{(i)}-c\textbf{I}\right)$
  for $k\leftarrow 2$ to $p$ do
     $\sigma_{2}\leftarrow\frac{1}{\gamma-\sigma}$
     $\textbf{R}_{\textbf{X}}\leftarrow\frac{2\sigma_{2}}{e}\textbf{A}\textbf{D}^{-1}\textbf{R}_{\textbf{Y}}-\frac{2\sigma_{2}}{e}c\textbf{R}_{\textbf{Y}}-\sigma\sigma_{2}\textbf{R}_{\textbf{X}}+\frac{2\sigma_{2}}{e}\textbf{Y}\boldsymbol{\Lambda}_{\textbf{Y}}$
     $\boldsymbol{\Lambda}_{\textbf{X}}\leftarrow\frac{2\sigma_{2}}{e}\boldsymbol{\Lambda}_{\textbf{Y}}\boldsymbol{\Lambda}^{(i)}-\frac{2\sigma_{2}}{e}c\boldsymbol{\Lambda}_{\textbf{Y}}-\sigma\sigma_{2}\boldsymbol{\Lambda}_{\textbf{X}}$
@@
  end for
  $\textbf{X}\leftarrow\textbf{D}^{-1}\textbf{R}_{\textbf{Y}}+\textbf{X}\boldsymbol{\Lambda}_{\textbf{Y}}$
@@ 4.2.1 Symmetric Eigenvalue Problems
From Fig. 7, we observe that for all the benchmark systems summarized in Table 1, the residual tolerance than can be achieved by the R-ChFSI method is nearly 10 orders of magnitude lower than what can be achieved using the ChFSI method when employing the diagonal approximation for the inverse overlap matrix.
@@
We also note that our proposed R-ChFSI algorithm allows us to construct the filtered subspace using lower precision and we report the residuals achieved by the FP32 and TF32 variants of R-ChFSI in Fig. 8. We also report the speedups achieved by the FP32 and TF32 variants of R-ChFSI over the FP64 variant of R-ChFSI in Fig. 9 on NVIDIA H100 GPU accelerators. We achieve speedups of up to 2.3x for the TF32 variant of the R-ChFSI method over the baseline implementation of the ChFSI method with FP64 arithmetic.
@@ 4.2.2 Hermitian Eigenvalue Problems
We now consider the finite-element discretization of the Kohn-Sham DFT equations at a non-zero $k$-point  [7] in the Brillouin zone for the systems described in Table 1. The resulting discretized equation is now a complex Hermitian eigenvalue problem. All the computations are done with FP64 according to the methodology prescribed by [7, 28].
@@
From Fig. 10, we observe that for all the benchmark systems summarized in Table 1, even for Hermitian eigenvalue problems the residual tolerance than can be achieved by the R-ChFSI method is nearly 10 orders of magnitude lower than what can be achieved using the ChFSI method when employing the diagonal approximation for the inverse overlap matrix.
@@
We also note that our proposed R-ChFSI algorithm allows for us to construct the filtered subspace using lower precision and we report the residuals achieved by the FP32 and TF32 variants of R-ChFSI in Fig. 11.
@@
We also report the speedups achieved by the FP32 and TF32 variants of R-ChFSI over the FP64 variant of R-ChFSI in Fig. 12 on NVIDIA H100 GPU accelerators. We achieve speedups of up to 2.7x for the TF32 variant of the R-ChFSI method over the baseline implementation of the R-ChFSI method with FP64 arithmetic.